OpenStudy (anonymous):

what is the derivative of y=x^(2/x), how to work it out? Thanks

OpenStudy (anonymous):

You use something called logarithmic differentiation: \[\eqalign{ &y=x^{\frac{2}{x}} \\ &\ln(y)=\ln(x^{\frac{2}{x}}) \\ &\ln(y)=\frac{2}{x}\ln(x) \\ &\ln(y)'=\left(\frac{2}{x}\ln(x)\right)' \\ &\frac{1}{y}y'=\left(\frac{2}{x^2}-\frac{2}{x^2}ln(x)\right) \\ &y'=y\left(\frac{2}{x^2}[1-ln(x)]\right) \\ &y'=\frac{2}{x^2}x^{\frac{2}{x}}[1-ln(x)] \\ &y'=2(x)^{\frac{2}{x}-2}[1-ln(x)] \\ }\]